OpenStudy (anonymous):

A triangular prism has a volume of 5 m3. If each dimension of the prism is tripled, what is the new volume of the prism? 45 m3 15 m3 135 m3 405 m3

OpenStudy (anonymous):

volume is proportional to the side cubed if sides are increased by three times, so, volume would increase by (3^3) =?

OpenStudy (anonymous):

Volume for triangular prism's should be V=1/2 * bhl=5. b= base length, h= height, l= length between 2 triangles. for the new volume, Lets say A. A=1/2 * (3b)(3h)(3l)=(1/2)bhl(3^3)=5*27=135. so i think the answer would be 13.

OpenStudy (anonymous):

@glitterMayhem 5*27 = 135m^3

OpenStudy (anonymous):

I don't understand this one bit -_-

OpenStudy (anonymous):

scaling a geometrical figure by "3" will 1) increase area by 3*3 = 9 2) increase volume by 3*3*3 = 27 in general, if a figure is scaled by some number "a", then 1) area is scaled by = a*a 2) volume is scaled by = a*a*a

OpenStudy (anonymous):

in the problem, the sides are tripled, that means, our prism is scaled by 3 so, the volume would scale by 3*3*3 = 27 so, the new volume is 27 times what it was before = 27 * 5 = 135m^3

OpenStudy (anonymous):

kapeesh?